OpenStudy (anonymous):

if the sum of a number and nine is doubled, the result is four less than the number. Find the number.

OpenStudy (midhun.madhu1987):

WELCOME TO OPENSTUDY Let the number be x. Then, (x+9) x 2 = x - 4 find x.

OpenStudy (brooke2353):

2(x+2)= 3x-6 2x+4= 3x-6 x=10